OpenStudy (anonymous):

.. What is the simplified form of 15 times x to the fifth power over 9 times y to the seventh power divided by 45 times y–cubed over 5 times x squared ? 5 x to the seventh over 27 y to the tenth power 15 x to the third power over y to the fourth y to the fourth power over 15 x–cubed 27 y to the seventh power over 5 x to the seventh power

OpenStudy (anonymous):

\[\Huge \frac{ \frac{15^5}{9y^7} }{ \frac{45y^3}{5x^2}}\]?

OpenStudy (anonymous):

yeahh..

OpenStudy (anonymous):

wow thats a biggy, ok lets work this out \[\Huge \frac{\frac{a}{b}}{\frac{c}{d}}=~\frac{a}{b}\times\frac{d}{c}\]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

So\[\Huge \frac{ \frac{15x^5}{9y^7} }{ \frac{45y^3}{5x^2}}=\frac{15x^{5}}{9y^{7}}\times\frac{5x^{2}}{45y^{3}}\]

OpenStudy (anonymous):

Is this a little easier to work with?